OpenStudy (anonymous):

1.Write and exponent . (-4)(-4) is it -4^2 i am so bad at this and i cannot wrap my head around it , can someone show me how to do it

OpenStudy (anonymous):

-4^2 just means (-4)(-4). If you had a problem that asked to write (-5)(-5)(-5)(-5) as a exponent, then your answer would be -5^how many (-5) there are...So its -5^4. Hope this helps :)

OpenStudy (anonymous):

Are you sure , its telling me to show my work ,

zepdrix (zepdrix):

\[\large -4^2 \ne (-4)^2\]Careful! Make sure to include the brackets! :)

OpenStudy (anonymous):

so the answer is (-4)2 :S?

zepdrix (zepdrix):

Mmmmm yah that's probably what you're suppose to do :) The wording on your question is kinda confusing lol. "Write and exponential" :o

zepdrix (zepdrix):

and exponent* blah...

OpenStudy (anonymous):

this is my question 1.write an exponent.show your work ,(-4)(-4)

OpenStudy (anonymous):

write using a exponent** sorry

zepdrix (zepdrix):

\[\large x \cdot x = x^2\]\[\large 3 \cdot 3 =3^2\] Applying this gives us, \[\large (-4)\cdot(-4)=(-4)^2\] So yah I guess that's what you were looking for :) To simply introduce an exponent so it looks simpler.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Yes (-4)(-4)=(-4)^2=16(positive) -(4)(4) =-4^2 =-16(negative)